# Can Automated Feedback Turn Students into Happy Prologians?

### v1

@@ 5.2 Impact of interest, engagement and LLM use (RQ2)
5.2.1 Interest
Figure 5: Answer distribution for the question “How interested are you in the topics of logic and logic programming?”. The median response is shown by the black vertical line.
@@ 5.2.1 Interest
Figure 6: Effect of interest in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
@@
5.2.2 Engagement
Figure 7: Effect of engagement in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
@@ 5.2.2 Engagement
5.2.3 Usage of LLMs
@@
Figure 8: Effect of LLMs usage in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
@@
Students were allowed to use Large Language Models (LLMs) during this course, as long as they disclosed that usage. Considering the 144 responses to our survey, 53 reported using LLMs when solving assignments, while 91 reported that they did not use LLMs. We were interested to know if having used LLMs has an effect on how students perceived the usefulness of ProHelp (for example, students who used LLMs might have found ProHelp less useful due to already having personalized feedback from the LLM). In Figure 8, we show the usefulness ratings for students who used and did not use LLMs. According to the Mann-Whitney $U$ test, there is no statistical difference between the two groups ($W=2261.50,\;p=0.51,\;r=-0.06$).
@@
There is no evidence that interest level, having used the platform for optional exercises, or having used LLMs, affects how helpful students found ProHelp overall.
5.3 Future direction (RQ3)
@@
Figure 9: Usefulness ratings for the different types of feedback not yet implemented in ProHelp. The median response for each type of feedback is shown by the black vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings.
@@
Our last goal is to better understand what features and improvements students would most like to have in the future, while also analyzing what aspects of the course students did not like. In the survey, we asked students to rate the 4 types of (not yet implemented) feedback described in Section 3.3 between Not Useful, Somewhat Useful and Useful. Figure 9 shows the students’ answers, where at least 75% of students rated each of the proposed feedback types as Useful. These responses hint that these are good goals to work towards.
@@
At the end of the survey, we asked students to suggest other types of feedback they would like to see in the future, as well as provide general comments/criticisms. One common criticism among students was that the unit tests were not specific enough, i.e., each test covered too much functionality. This left students with little information when tests failed and highlights the importance of faculty creating good unit tests that assess each functionality in isolation. There were also other comments regarding the course organization and grading method, which we have relayed to the faculty.
@@
Although students found all proposed types of feedback useful, there is a significant preference for the “Show the outputs and expected outputs” compared with the remaining feedback types. Students would like to have an automated estimation for the qualitative component of their projects.
6 Discussion and Conclusion
This work proposes ProHelp, a new tool that provides automated feedback for students learning logic programming using Prolog. Additionally, we explore more advanced types of automated feedback that have not yet been implemented. Our main contribution is a student survey with 144 valid responses, gathering insights about students’ preferences and complaints regarding automated assessment tools, specifically GitSEED and ProHelp.
Survey analysis indicates that users found the feedback received useful and the GitSEED interface was user-friendly. The components of ProHelp that students most appreciated include automatic testing, warnings about open choice points, and scoring for each required predicate. Students also displayed great interest in the more advanced types of feedback we proposed and suggested some themselves. These findings will be valuable for future research, enabling educators and developers to focus on the most impactful types of feedback.
@@ 6 Discussion and Conclusion
Building on the results of the survey, we plan to enhance ProHelp with the following improvements: (1) use fault localization and program diagnosis techniques to give more information to students when tests fail, particularly for failed assertions and infinite loops, (2) explore automated repair methods, namely program synthesis and LLMs, to generate correction hints or bug explanations for students, and (3) investigate LLM-based techniques to estimate qualitative aspects of logic programs. One challenge highlighted by some students is that excessive feedback may hinder learning rather than help. Balancing the amount and type of feedback will be crucial to ensuring that students receive meaningful guidance without becoming overly reliant on automation.
In conclusion, this work introduces ProHelp, a tool offering basic automated feedback for Prolog, and presents findings from a user study on students’ preferences for automated assessment tools. The results indicate that students found ProHelp valuable, particularly its automatic testing and reporting features, and showed strong interest in future enhancements.

### v2

@@ 6.2. Impact of interest, engagement and LLM use (RQ2)
6.2.1. Interest
Figure 6. Answer distribution for the question “How interested are you in the topics of logic and logic programming?”. The median response is shown by the black vertical line.
@@ 6.2.1. Interest
Figure 7. Effect of interest in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
@@
6.2.2. Engagement
Figure 8. Effect of engagement in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
@@ 6.2.2. Engagement
6.2.3. Usage of LLMs
Table 1. Number of students that reported using each LLM tool.
LLM Tool Count
ChatGPT 82
Copilot 16
Claude 5
Gemini 2
Blackbox AI 1
Deepseek 1
Perplexity AI 1
ZZZ Code AI 1
@@ 6.2.3. Usage of LLMs
Figure 9. Effect of LLMs usage in the agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
@@
Students were allowed to use Large Language Models (LLMs) during this course, as long as they disclosed that usage. Considering the 144 responses to our survey, 91 reported using LLMs when solving assignments, while 53 reported that they did not use LLMs. Table 1 shows the number of students that reported using each LLM tool/platform, with the most popular ones being ChatGPT and Copilot. Furthermore, 21 students reported using more than one LLM. We were interested to know if having used LLMs has an effect on how students perceived the usefulness of ProHelp (for example, students who used LLMs might have found ProHelp less useful due to already having personalized feedback from the LLM). In Figure 9, we show the usefulness ratings for students who used and did not use LLMs. According to the Mann-Whitney $U$ test, there is no statistical difference between the two groups (W=2261.50,p=0.51,r=0.06W=2261.50,\;p=0.51,\;r=-0.06).
@@
A2. There is no evidence that interest level, having used the platform for optional exercises, or having used LLMs, affects how helpful students found ProHelp overall.
6.3. Submission Analysis (RQ3)
Table 2. Number of correct and incorrect submission per assignment.
Assignment Correct Incorrect Total Avg. Clauses
Role-playing Exercise 271 239 510 3.51
    Day 1 88 58 146 3.00
    Day 2 63 67 130 3.27
    Day 3 40 68 108 5.05
    Day 4 46 9 55 2.62
    Day 5 34 37 71 3.26
Recitation Exercises 969 442 1411 2.50
    Exercise A1 65 55 120 2.00
    Exercise A2 79 44 123 2.67
    Exercise A3 61 24 85 2.05
    Exercise A4 49 15 64 2.00
    Exercise A5 49 12 61 3.05
    Exercise A6 50 21 71 2.00
    Exercise A7 46 19 65 2.08
    Exercise B1 57 42 99 3.09
    Exercise B2 39 31 70 1.77
    Exercise B3 42 16 58 3.55
    Exercise B4 41 7 48 1.70
    Exercise B6 45 15 60 2.86
    Exercise C1 41 19 60 2.22
    Exercise C2 36 6 42 2.98
    Exercise C3 35 7 42 3.32
    Exercise C4 37 17 54 3.11
    Exercise D1 47 13 60 2.25
    Exercise D2 35 10 45 2.56
    Exercise D3 36 14 50 3.09
    Exercise D4 33 23 56 3.02
    Exercise D5 23 27 50 2.12
    Exercise D6 23 5 28 2.00
Project 440 4840 5280 34.20
Total 1680 5521 7201
Figure 10. Number of submissions over time for each type of assignment.
Of the 312 students that interacted with ProHelp, 265 gave permission for their code to be used for research purposes. Over the course, those 265 students created a total of 7201 code submissions to the different repositories. We have created a publicly available annotated dataset containing these submissions (Brancas et al., 2025b). Table 2 shows the total number of correct and incorrect submissions per assignment, along with the average number of clauses in the students’ submissions for each one. Overall, there were 1680 submissions that passed all the tests for the respective assignment, and 5521 submissions that failed at least one test. By far, the assignment with the most number of submissions is the final project, which is easily explained by it being the only mandatory and graded assignment. The ratio of correct to incorrect submissions also varies greatly, with the optional assignments having much fewer incorrect submissions than the project. There are two plausible explanations for this: (1) the optional assignments are easier, and thus take fewer tries to get right, and (2) only more engaged students solved the optional exercises. Figure 10 shows the number of submissions for each day in the course and allows us to spot some interesting patterns in students’ behavior. Particularly, that most students only start solving the optional exercises on the platform once the project submissions are opened. It is also possible to see the escalating number of submissions on the three days leading to the project deadline.
Figure 11. Classification of the 7201 student submissions.
We analyzed the 7201 submissions and classified them along two dimensions: (1) is the submission correct or incorrect, and (2) is the submission an improvement or a regression compared with the previous version of the code. For this second category we consider 5 different labels:
  • First Submission: first submission of a student for a particular assignment;
  • Bug Fixed: the submission passes strictly more tests than the previous version;
  • Bug Introduced: the submission passes strictly fewer tests than the previous version;
  • Mixed: the submission both fails tests that previously passed and passes tests that previously failed;
  • No Change: the submission passes exactly the same tests as the previous version.
Figure 11 shows the number of submissions in each category. We observe that most of the first submissions are correct, but, as evidenced by Table 2, this is mostly due to the optional exercises. We also observe that the number of regressions is fairly small (Mixed, Bug Introduced). Lastly, there are a small number of submissions that are purely code quality improvements/refactorings of submissions that were already correct.
In order to understand the types and frequency of bugs committed by students, we analyzed the submissions in the Bug Fixed label. Of those 2495 submissions, we isolated the ones where only one predicate was modified – this subset represents “ideal” unit commits where students fixed a specific incorrect behavior in their program. This selection resulted in 767 submissions, of which we manually classified 200 chosen randomly.
LABEL:tab:bug-types presents the identified bug types and subtypes found in the 200 selected programs, along with their frequencies and a representative example for each type. Note that a single program may contain multiple bugs, and thus the total number of reported bug types exceeds 200. The most commonly identified reason for failing submissions was “Incomplete” programs, which can happen for two reasons: (1) students implemented their programs incrementally and tested components as they progressed (most often represented by the “Missing Predicate” subtype), and (2) students forgot to implement some functionality (most often represented by the “Missing Clause” subtype). While the first case does not necessarily represent a true bug, it still gives us the relevant insight that students seem to like having many tests in order to evaluate each functionality in isolation.
Other bug types were more evenly distributed, with no specific bug occurring at an overwhelmingly high frequency. Of particular note is the “Domain Logic Problem” bug type, which reflects a misunderstanding of the intended functionality for a given predicate. Additionally, 13 of the 26 programs categorized as “Other” contain at least one predicate that is fundamentally incorrect and requires complete rewriting.
A3. We have identified 9 categories of bug types committed by students, with 17 total subtypes. We also created a public dataset with 7201 Prolog programs, including 200 manually annotated with bug types and corrections.
Table 3. Description and frequency of the different bug types identified in student submissions.
Bug Type Count Frequency
Incomplete 75 37.5%
  Subtypes
     Missing Predicate 59 29.5%
     Missing Clause (of existing predicate) 16 8.0%
There are missing predicates/clauses. Example: +mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N, mult(L1, N, L2).
Wrong Argument 54 27.0%
  Subtypes
     Argument Order Swap 7 3.5%
     Missing Argument 6 3.0%
     Extra Argument 2 1.0%
     Other 39 19.5%
Issue in the arguments of a predicate call. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N, mult(L1, N, [-L1-]{+L2+}).
Rule Goal Problems 26 13.0%
  Subtypes
     Extra Goal 13 6.5%
     Missing Goal 10 5.0%
     Goal Order Swap 3 1.5%
Problems with the goals in the body of a rule. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N{+, mult(L1, N, L2)+}.
Operator Error 23 11.5%
  Subtypes
     Wrong Operator 16 8.0%
     List Terminators Issue 4 2.0%
     Missing Negation 3 1.5%
Incorrect operator usage. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 [-=-]{+is+} E1 * N, mult(L1, N, L2).
Wrong Predicate Name 19 9.5%
Incorrect predicate name in call. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * N, mult[-iply-](L1, N, L2).
Domain Logic Problem 16 8.0%
Different problems on the implementation of domain logic.
Cut Problem 14 7.0%
  Subtypes
     Missing Cut 10 5.0%
     Extra Cut 3 1.5%
     Wrong Placement 1 0.5%
Problems with the cut operator (!). Example: max(X, Y, X) :- X >= Y{+, !+}. max(_, Y, Y).
Wrong Variable/Constant 8 4.0%
  Subtypes
     Wrong Variable Name 6 3.0%
     Wrong Constant 2 1.0%
Problems with the names of variables or constants. Example: mult([], _, []). mult([E1 | L1], N, [E2 | L2]) :- E2 is E1 * [-M-]{+N+}, mult(L1, N, L2).
Other 26 13.0%
Table 3. (continued)
@@ 6.3. Submission Analysis (RQ3)
Figure 12. Usefulness ratings for the different types of feedback not yet implemented in ProHelp. The median response for each type of feedback is shown by the black vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings.
6.4. Future directions (RQ4)
@@
Our last goal is to better understand what features and improvements to the feedback students would most like to have in the future, while also analyzing what aspects of the course students did not like. In the survey, we asked students to rate the 4 types of (not yet implemented) feedback described in Section 4.2 between Not Useful, Somewhat Useful and Useful. Figure 12 shows the students’ answers, where at least 75% of students rated each of the proposed feedback types as Useful. These responses hint that these are good goals to work towards.
@@
Table 4. Description of issues and features raised by students in open-ended responses, along with the count of how many students voiced each one. Issues/features reported by only one student were omitted.
Description Count
Predicate scores should be proportional to the number of tests passed 11
Tests should be more specific and in greater quantity 10
The project statement and/or organization of the repository was confusing 5
It should be clear what predicates each test is targeting 4
Having more feedback might not lead to better learning 3
Instructions for how to use Git/VS Code/etc were confusing 3
The faculty was not very responsive on Discord 3
Feature: show test execution times 2
Feature: predict the qualitative grade portion of the evaluation 2
Waiting 15 minutes between submissions is too long 2
The way submissions are counted was unfair 2
At the end of the survey, we asked students to suggest other types of feedback they would like to see in the future, as well as provide general comments/criticisms. Table 4 shows the issues raised by students. One common criticism among students was that the unit tests were not specific enough, i.e., each test covered too much functionality. For example, P-31 said: “The project tests were not exhaustive enough.”. This left students with little information when tests failed and highlights the importance of faculty creating good unit tests that assess each functionality in isolation. There were also other comments regarding the course organization and grading method, which we have relayed to the faculty.
@@
A4. Although students found all proposed types of feedback useful, there is a significant preference for the “Show the outputs and expected outputs” compared with the remaining feedback types. Students would like to have an automated estimation for the qualitative component of their projects.
7. Discussion and Conclusion
This work analyses students’ behaviors during a 9-week long undergraduate class on logic programming. Specifically, we study how they interacted with an automated assessment platform for Prolog code. This platform, named ProHelp, included several types of automated feedback to guide students to find and solve their bugs more easily. Our main contributions are: the analysis of a student survey with 144 valid responses, gathering insights about students’ preferences and complaints regarding ProHelp, and a dataset of 7201 Prolog programs with 200 manually annotated ones with bug types and corrections.
Survey analysis indicates that users found the feedback received useful and the ProHelp interface was user-friendly. For instance, P-101 said “I loved the assessment system. It was very efficient and helpful. From what I have heard from previous years, it was a huge improvement”. The components of ProHelp that students most appreciated include automatic testing, warnings about open choice points, and scoring for each required predicate. Students also displayed great interest in the more advanced types of feedback we proposed and suggested some themselves. These findings will be valuable for future research, enabling educators and developers to focus on the most impactful types of feedback.
@@ 7. Discussion and Conclusion
Our submission code analysis resulted in a public dataset which can be used by future work to develop and test fault localization and program repair techniques. The report on the types of bugs students commit most frequently can also be used by faculty in order to stress those point a priori in-class. Likewise, insights such as students only starting to solve optional exercises once that knowledge is also required for graded work can also be used by faculty in the development of future courses.
@@
Building on the results of the survey and the submissions dataset, we plan to enhance ProHelp with the following improvements: (1) use fault localization and program diagnosis techniques to give more information to students when tests fail, particularly for failed assertions and infinite loops, (2) explore automated repair methods, namely program synthesis and LLMs, to generate correction hints or bug explanations for students, and (3) investigate LLM-based techniques to estimate qualitative aspects of logic programs. One challenge highlighted by some students is that excessive feedback may hinder learning rather than help. Balancing the amount and type of feedback will be crucial to ensuring that students receive meaningful guidance without becoming overly reliant on automation.
Acknowledgements.
This work was partially supported by Portuguese national funds through FCT, under project 2023.14280.PEX (DOI: 10.54499/2023.14280.PEX), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (FUN2MODEL, grant agreement No. 834115). This work was also supported by the Carnegie Mellon University Portugal Program and FCT under grants PRT/BD/152086/2021 (DOI: 10.54499/PRT/BD/152086/2021) and PRT/BD/153739/2021 (DOI: 10.54499/PRT/BD/153739/2021).
Appendix A Survey Protocol
The following survey was designed to collect anonymous feedback from undegraduate students enrolled in the Logic Programming course regarding their experience with the automated feedback system provided via GitLab. The survey was estimated to take approximately 10 to 15 minutes. Several questions were included to assess usability, usefulness, and pedagogical effectiveness of the feedback tools. Some questions functioned as attention checks. The questions and answers presented here were translated from Portuguese.
Section 1: Demographics and Background
  • What gender do you identify with?
    • Female
    • Male
    • Prefer not to say
    • Prefer to self-describe
  • What is your level of interest in the topics covered in the Programming Languages course? (Logic, Prolog)
    Rating scale from 1 (not interested) to 5 (very interested)
  • During this course, did you experience difficulties using Git/GitLab? Select all that apply:
    • Merge conflicts
    • SSH key problems
    • Problems interacting with VS Code
    • I had no problems
    • Other
Section 2: Automated Feedback in Prolog
Students were reminded of the types of automated feedback they encountered on GitLab:
  • Automatic test verification
  • Predicate scoring
  • Validation of solution types
  • Dashboards with scores
  • Syntax error highlighting
  • Suggestions for incorrect predicate names
  • Warnings about open choice points
Students were asked to rate their agreement with the following System Usability Survey Items using a 5-point Likert scale (1 = Strongly Disagree, 5 = Strongly Agree):
  • I think that I would like to use this system frequently.
  • I found the system unnecessarily complex.
  • I thought the system was easy to use.
  • I think that I would need the support of a technical person to be able to use this system.
  • I found the various functions in this system were well integrated.
  • I thought there was too much inconsistency in this system.
  • I would imagine that most people would learn to use this system very quickly.
  • I found the system very cumbersome to use.
  • I felt very confident using the system.
  • I needed to learn a lot of things before I could get going with this system.
Additionally, students rated the usefulness of each feedback component on a scale from 1 (Not useful at all) to 5 (Very useful):
  • Suggestions for incorrect predicate names
  • Syntax error highlighting
  • Dashboards with scores
  • Select “Not useful at all” in this line [attention check]
  • Automatic test verification
  • Predicate scoring
  • Warning for open choice points
  • Solution type validation (iterative, recursive, functional)
Section 4: Role-playing Exercise
  • Did you solve the role-playing exercise?
    • No
    • I started but did not finish
    • Yes
  • Why did you not solve/abandon/finish the exercises? (Open text response)
  • Did you use LLMs (e.g., ChatGPT, Copilot) to help solve the role-playing exercise?
    • No
    • Yes
  • Which LLM(s) did you use? (Open text response)
  • What did you think of the sequence of puzzles from the role-playing exercise on a 5-point scale (1 = Not entertaining at all, 5 = Very entertaining)?
Section 5: Recitation Exercises
  • Did you solve the recitation exercises using Git/GitLab?
    • No
    • I started but did not finish
    • Yes
  • Why did you not solve/abandon/finish the exercises? (Open text response)
  • Did you use LLMs (e.g., ChatGPT, Copilot) to help with the recitation exercises?
    • No
    • Yes
  • Which LLM(s) did you use? (Open text response)
Section 6: Project
  • Did you use LLMs (e.g., ChatGPT, Copilot) to help with the project?
    • No
    • Yes
  • Which LLM(s) did you use? (Open text response)
Section 7: Final Reflections
  • Please rate the following statements on a 5-point Likert scale (1 = Strongly Disagree, 5 = Strongly Agree)
    • I believe the automated feedback I received on GitLab improved my grade in this course.
    • I believe the role-playing exercise improved my grade in this course.
    • Select “Strongly Agree” to demonstrate you are paying attention. [Attention check]
    • I believe the recitation exercises on GitLab improved my grade in this course.
  • Please rate the potential utility of proposed features on a 3-point scale (Indifferent, Somewhat Useful, Useful):
    • Suggestions for corrections on incorrect submissions
    • Showing program output for each test compared to expected output
    • Automatic identification of bug locations
    • Detection of infinite loops
  • Please suggest any other features you would like to see. (Open text response)
  • Please leave any other feedback you may have (positive or negative). (Open text response)